# Periodic orbits and their gravitational waves in EMRIs: supermassive black hole affected by galactic dark matter halos

### v1

@@ Abstract
Periodic orbits exhibiting zoom-whirl behavior have become attractive topics for studying particle dynamics and gravitational wave emission in extreme-mass-ratio inspirals (EMRIs). This study systematically investigates periodic orbits around black holes and their gravitational wave radiation in three dark matter halo environments: NFW, Beta, and Moore models. The dark matter distribution in these models can be effectively incorporated using two parameters—the dark matter characteristic mass and halo characteristic radius. Our results reveal that for a larger dark matter mass and a smaller characteristic radius, the shapes of the periodic orbits and the corresponding gravitational waveforms show more significant deviations from the Schwarzschild case. As the halo characteristic radius increases, the orbital shapes and waveform characteristics gradually converge with the Schwarzschild black hole results. For the different dark matter halo models, the NFW and Beta models produce nearly indistinguishable results, while the Moore model shows distinctly different effects. These findings enhance our understanding of dark matter halo effects on gravitational wave signatures and may provide constraints for future space-based detectors.
@@ I Introduction
The motion of test particles (or small celestial bodies) in a gravitational field serves as a pivotal approach to probe the properties of gravitational fields. Bound precession orbits and circular orbits play a foundational role in the classical tests of gravitational theories, the long-term stability and evolution of gravitational systems (especially strong-field regimes), and gravitational wave emission [29, 30, 31, 32, 33, 34, 35, 36, 37]. However, the motion of small celestial bodies and test particles in strong gravitational fields may also exhibit another important motion pattern—periodic orbits [38, 39, 40, 41]. Periodic orbits have gradually become an inspiring research topic in recent years. They have been extensively studied in quantum-corrected spacetime under quantum gravity theory [42, 43, 44, 45, 46, 47], hairy black holes [48, 49, 50], and other spacetime backgrounds [51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. The periodic orbits universally exist in various gravitational systems, and their configuration is typically determined by three integer values: the zoom number $z$, the whirl number $w$, and the vertex number $v$ [38, 39, 41]. Specifically, $z$ denotes the number of “leaves” in a full periodic cycle, with greater $z$ values indicating increased trajectory complexity; $w$ represents the count of additional whirls the particle performs during its outward drift toward the apoapsis; and $v$ characterizes the behavior of the subsequent vertex the particle encounters after leaving the initial vertex (apoapsis). Moreover, the zoom and whirl features of periodic orbits are directly reflected in the gravitational waveform. These periodic orbits and their gravitational-wave radiation signals are promising probes for the exploration of strong gravitational fields. These signals are particularly important for the study of extreme-mass-ratio inspirals (EMRIs) [63, 43, 64]. EMRIs are of significant importance in gravitational wave astronomy. They are composed of a compact small celestial body/test particle that gradually spirals towards the SMBH [65]. They have been identified as promising targets for the next generation of gravitational-wave detectors, such as LISA [66, 67], TianQin [68, 69], Taiji [70, 71] and DECIGO [72].
@@
Recent theoretical studies have examined the influence of dark matter on the physical processes in the vicinity of SMBHs. These studies have involved a range of topics, including accretion [73, 74], gravitational lensing [75, 76, 77, 78, 79, 80, 81], circular geodesics [82, 83, 84] and quasi-normal mode and gravitational waves [85, 86, 87, 88]. However, the studies on the periodic orbits around SMBH in dark matter environments remain in a very preliminary stage [89, 64]. The specific impact of dark matter mass and density profiles in halo structures on the periodic orbit of particles around SMBHs, as well as the subsequent influence on the generated gravitational wave signals, remain an unresolved and challenging issue. These questions are of significant importance for the interpretation of gravitational wave signals in EMRIs in gravitational wave astronomy.
@@
Based on the aforementioned motivations, this study focuses on the properties of periodic orbits around supermassive black holes in dark matter halo environments. Specifically, we select three typical dark matter halo models (NFW, Beta, and Moore), whose applicability has been tested by theoretical investigations and astrophysical observations in galaxies. The trajectory characteristics of test particle periodic orbits under different dark matter halo scales $h$ and dark matter masses $k$ are analyzed through numerical simulations. Meanwhile, we systematically investigate the gravitational wave signals generated by periodic orbits in EMRIs influenced by these dark matter models, quantifying the impacts of different dark matter halos on the shape of periodic orbits and gravitational waveforms. The analysis in this paper is crucial for gaining a deeper understanding of how different dark matter density distributions precisely affect particle orbits around SMBHs, and it may provide value for capturing observable signal features in future space-based gravitational wave detections.
@@ II.2 Effective Potential, ISCO and MBO
Figure 1: The influence of dark matter on the effective potential of particles. (a) The left subplot highlights the effects of different dark matter halo models, with $k=1\times 10^{4}M$. (b) The right subplot emphasizes the role of the dark matter mass $k$. In both subplots, the particle angular momentum is set to $L=3.8$, and the dark matter halo characteristic radius is $h=1\times 10^{7}M$. The black curve corresponds to the case without dark matter (i.e., the Schwarzschild black hole).
@@ II.3 Periodic Orbits and Their Precession Parameter $q$
$q=\frac{\omega_{\phi}}{\omega_{r}}-1=\frac{\Delta\phi}{2\pi}=w+\frac{v}{z}.$ (17)
@@ IV.2 The effect of dark matter halo scale on gravitational waves
Following the analysis of dark matter mass effects, we now examine how the dark matter halo scale $h$ influences gravitational wave signals. Fig. 17 presents the waveforms produced by the (221)(2~2~1) orbital configuration for varying dark matter halo scales in the three models. The dark matter mass is maintained constant as $k=10^{4}M$, while the halo characteristic radius $h$ varies from $10^{7}M\sim 10^{10}M$. The several dark matter models produce distinguishable features in gravitational waveforms for a relatively small dark matter halo scale, and the results reveal a convergence pattern as the halo characteristic radius increases. When $h=10^{7}M$, the gravitational waveforms obtained in three dark matter models exhibit noticeable deviations from the Schwarzschild reference baseline, reflecting the significant influence of the dark matter halo on the spacetime geometry and particle orbits. These deviations are evident in multiple waveform characteristics: the signal period extends to approximately $1200M$ in a time period (compared to roughly $950M$ for Schwarzschild results), and the oscillation period is noticeably longer. As the halo characteristic radius expands to $h=10^{8}M$, these deviations diminish substantially, with the waveforms moving closer to the Schwarzschild case. The signal duration contracts toward the Schwarzschild value, and the oscillation periods become more similar. This convergence process continues as $h$ increases further to $h=10^{9}M$, at which point the waveforms of all dark matter models become nearly indistinguishable from the Schwarzschild reference baseline in terms of signal duration, oscillation period and amplitude structure. When $h>10^{9}M$, the periodic orbits of the dark matter halo models completely coincide with those of the Schwarzschild black hole, and the corresponding gravitational waveforms are identical as well. For this reason, the results for such cases are not presented separately. This convergence behavior directly parallels the halo effects observed in the precession angle analysis (Figs. 8 and 9) and orbital trajectories (Fig. 10). The physical interpretation remains consistent: when the dark matter halo characteristic radius becomes sufficiently large while maintaining dark matter total mass as constant, the dark matter halo is diluted and the local gravitational field gradually approaches that of an isolated Schwarzschild black hole. The diluted halo configuration results in weaker gravitational binding in the regions where periodic orbits are located, thereby reducing the distinguishable signatures of different dark matter density profiles in the gravitational waves.
@@
To provide a more comprehensive view of dark matter halo’s scale effects across different orbital configurations, Fig. 19 presents gravitational waveform comparisons for varying NFW halo scales using five orbital configurations: (110)(1~1~0), (120)(1~2~0), (130)(1~3~0), (211)(2~1~1), and (221)(2~2~1). Four scale values are examined: $h=10^{7}M$, $h=10^{8}M$, $h=10^{9}M$ and $h=10^{10}M$, with the Schwarzschild case serving as the reference baseline. These results not only reveal how different orbital configurations (zwv)(z~w~v) modulate the gravitational waveforms (consistent with findings in previous sections), but also indicate the clear trend: larger dark matter halo scales lead to weaker effects on the gravitational potential, with waveforms approaching the Schwarzschild black hole case.
@@ V Conclusion
Our analysis of the effective potential reveals that the introduction of dark matter halos reduces the extrema of the effective potential. This reduction becomes more obvious as the dark matter mass parameter $k$ increases, directly affecting the MBO and ISCO of particles. The presence of dark matter halos increases the angular momentum for both MBO and ISCO, with the NFW and Beta models showing nearly identical behavior for the parameter ranges considered, while the Moore model exhibits systematically lower angular momentum requirements.
@@
For periodic orbits, we examine the effects of both dark matter mass $k$ and characteristic radius $h$ under the condition of selected angular momentum $L=L_{\text{ISCO}}+\varepsilon(L_{\text{MBO}}-L_{\text{ISCO}})$ and the selected energy $E$ (in Appendix B). For the selected angular momentum condition, we find that for a given precession angle $q$, larger dark matter masses require lower orbital energies $E$. The NFW and Beta models require substantially lower energies $E_{(z~w~v)}$ than the Moore model for the same precession angle and orbital configurations (zwv)(z~w~v). This energy difference directly affects the orbital shapes increasing dark matter mass stretches the apoapsis of periodic orbits, leading to larger orbital dimensions. The scale dependence presents a complementary pattern. For a fixed precession angle $q$, larger dark matter halo scales result in higher required orbital energies. At smaller scales, orbital shapes differ from the Schwarzschild case, but as the scale increases, the orbits gradually converge toward the Schwarzschild trajectories, closely approaching them at large scales. This behavior reflects the dilution of gravitational effects: when the dark matter mass remains constant while the scale increases, the gravitational binding weakens in the regions. Particles must then possess higher energies to maintain the same precession characteristics. This scale-dependent effect is clearly visible in the periodic orbit configurations shown in Fig. 10, where orbits gradually approach the Schwarzschild case as $h$ increases. In particular, when the dark matter halo scale is $h\sim 10^{10}M$, the periodic orbits of all three dark matter models become indistinguishable from the Schwarzschild reference, demonstrating the dilution limit of dark matter halo effects. Furthermore, analysis with fixed $E$ yields conclusions consistent with those obtained for fixed $L$.
@@
We analyze an extreme-mass-ratio inspiral (EMRI) system consisting of a test object with a typical mass ratio and luminosity distance. The gravitational wave analysis shows strong correspondence between the waveforms and the orbital properties. Larger dark matter mass parameters lead to more pronounced differences in generated gravitational waves between black holes in dark matter halos and those of Schwarzschild black holes, with enhanced distinguishability among the three dark matter halo models. The waveforms exhibit stretching, particularly an increase in the duration of each complete orbital period, which stems directly from the modification of the gravitational potential by the dark matter halo. Examining the waveform structure reveals that the orbital parameters (zwv)(z~w~v), as fundamental quantities characterizing the geometric shapes of periodic motion, directly determine the spatiotemporal features of the gravitational wave signals.The dark matter mass $k$ influences the waveforms indirectly by modifying the effective gravitational potential, thereby extending orbital periods and broadening the characteristic timescales of the gravitational wave signals. In contrast to the mass effects, a larger dark matter halo scale $h$ results in a more diffuse mass distribution, weakening its effect on the gravitational potential. The gravitational waveforms generated by periodic orbits in dark matter halo models gradually approach those for Schwarzschild black holes when dark matter halo scales become sufficiently large. The results presented in this work indicate that larger dark matter halo scale $h$ corresponds to weaker modification of the gravitational potential, leading to waveforms that more closely resemble the Schwarzschild case. Particularly, when the characteristic radius of the dark matter halo is $h\sim 10^{10}M$, the periodic orbits and their gravitational waveforms converge precisely with those of the Schwarzschild black hole.
@@
Throughout our analysis, we observe that the NFW and Beta models produce nearly indistinguishable results across both orbital characteristics and gravitational wave signals for all parameter ranges examined in this study. This equivalence holds for precession angles, periodic orbit configurations, and gravitational waveforms, suggesting that these two density profiles have very similar effects on particle orbits and gravitational radiation in the mass and radius ranges relevant to galactic dark matter halos. The Moore model, while following similar qualitative trends, generates systematically different quantitative results that become more obvious at larger dark matter masses.
@@
In summary, through systematic analysis of dark matter halo models, this study clarifies how dark matter halo parameters influence particle orbits around black holes and gravitational wave radiation. The complementary effects of the dark matter mass and its characteristic radius offer a deeper understanding of how dark matter affects particle orbits, as well as provide theoretical predictions on gravitational waveforms that can be compared with observations. These findings may provide useful constraints for interpreting gravitational wave signals from EMRIs detected by future space-based observatories, such as LISA, TianQin, and Taiji.

### v2

@@ Abstract
Periodic orbits exhibiting zoom-whirl behavior have become attractive topics for studying particle dynamics and gravitational wave emission in extreme-mass-ratio inspirals (EMRIs). This study systematically investigates periodic orbits around black holes and their gravitational wave radiation in three dark matter halo environments: NFW, Beta, and Moore models. The dark matter distribution in these models can be effectively incorporated using two parameters—the dark matter characteristic mass and halo characteristic radius. Our results reveal that for a larger dark matter mass and a smaller characteristic radius, the shapes of the periodic orbits and the corresponding gravitational waveforms show more significant deviations from the Schwarzschild case. As the halo characteristic radius increases, the orbital shapes and waveform characteristics gradually converge with the Schwarzschild black hole results. Furthermore, our results also suggest that the NFW and Beta models produce nearly indistinguishable results, while the Moore model shows distinct and resolvable signatures compared with Beta/NFW models. These findings deepen our understanding of dark matter halo effects on periodic motions and gravitational wave signatures, providing guidance on future space-based observation for EMRIs.
@@ I Introduction
The motion of test particles (or small celestial bodies) in a gravitational field serves as a pivotal approach to probe the properties of gravitational fields. Bound precession orbits and circular orbits play a foundational role in the classical tests of gravitational theories, the long-term stability and evolution of gravitational systems (especially strong-field regimes), and gravitational wave emission [29, 30, 31, 32, 33, 34, 35, 36, 37]. However, the motion of small celestial bodies and test particles in strong gravitational fields may also exhibit another important motion pattern—periodic orbits [38, 39, 40, 41]. Periodic orbits have gradually become an inspiring research topic in recent years. They have been extensively studied in quantum-corrected spacetime under quantum gravity theory [42, 43, 44, 45, 46, 47], hairy black holes [48, 49, 50, 51], and other spacetime backgrounds [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66]. The periodic orbits universally exist in various gravitational systems, and their configuration is typically determined by three integer values: the zoom number $z$, the whirl number $w$, and the vertex number $v$ [38, 39, 41]. Specifically, $z$ denotes the number of “leaves” in a full periodic cycle, with greater $z$ values indicating increased trajectory complexity; $w$ represents the count of additional whirls the particle performs during its outward drift toward the apoapsis; and $v$ characterizes the behavior of the subsequent vertex the particle encounters after leaving the initial vertex (apoapsis). Moreover, the zoom and whirl features of periodic orbits are directly reflected in the gravitational waveform. These periodic orbits and their gravitational-wave radiation signals are promising probes for the exploration of strong gravitational fields. These signals are particularly important for the study of extreme-mass-ratio inspirals (EMRIs) [67, 43, 66]. EMRIs are of significant importance in gravitational wave astronomy. They are composed of a compact small celestial body/test particle that gradually spirals towards the SMBH [68]. They have been identified as promising targets for the next generation of gravitational-wave detectors, such as LISA [69, 70], TianQin [71, 72], Taiji [73, 74] and DECIGO [75].
@@
Recently, theoretical studies have examined the influences of dark matter on the physical processes in the vicinity of SMBHs. These studies involve a wide range of topics and physical processes, including accretion [76, 77], gravitational lensing [78, 79, 80, 81, 82, 83, 84], circular geodesics and black hole shadow  [85, 86, 87, 88, 89, 90, 91, 92], quasi-normal mode and gravitational waves [93, 94, 95, 96, 97, 98]. However, the studies on periodic orbits around SMBH in dark matter environments remain in a very preliminary stage, which are concentrated on a single dark matter model [65, 66]. The comparison of periodic orbits from different dark matter distributions is still absent. The specific impact of dark matter mass and density profiles in halo structures (described by different halo models) on the periodic orbits around SMBHs, as well as the subsequent influence on the generated gravitational wave signals, remain unresolved and challenging issues. These questions are of significant importance for the interpretation of gravitational wave signals in EMRIs in gravitational wave astronomy.
@@
Inspired from the aforementioned motivations, this study focuses on the properties of periodic orbits around supermassive black holes in dark matter halo environments. This work enables us to give a comprehensive analysis on dark matter influences from different halo models. Specifically, we select three typical dark matter halo models (NFW, Beta, and Moore), whose validity has been tested by theoretical investigations and astrophysical observations in galaxies [99]. The characteristics of test particles’ periodic orbits calculated within these halo models under different dark matter masses $k$ and halo scales $h$ are analyzed through numerical calculations. Meanwhile, we investigate the gravitational wave signals generated by periodic orbits in EMRIs influenced by these dark matter models, clarifying the impacts of different dark matter halos on the shape of periodic orbits and gravitational waveforms. The analysis presented in this work would be helpful for gaining a deeper understanding of how different dark matter density distributions precisely affect particle orbits around SMBHs, and it may provide potential applications for capturing observable signatures in future space-based gravitational wave detections.
@@ II.2 Effective Potential, ISCO and MBO
Figure 1: The influence of dark matter on the effective potential of particles. (a) The left subplot highlights the effects of different dark matter halo models, with $k=10^{4}M$. (b) The right subplot emphasizes the role of the dark matter mass $k$. In both subplots, the particle angular momentum is set to $L=3.8$, and the dark matter halo characteristic radius is $h=10^{7}M$. The black curve corresponds to the case without dark matter (i.e., the Schwarzschild black hole).
@@ II.3 Periodic Orbits and Their Precession Parameter $q$
$q=\frac{\Delta\phi}{2\pi}-1=w+\frac{v}{z}.$ (17)
@@ IV.2 The effect of dark matter halo scale on gravitational waves
Following the analysis of dark matter mass effects, we now examine how the dark matter halo scale $h$ influences gravitational wave signals. Fig. 17 presents the waveforms produced by the (221)(2~2~1) orbital configuration for varying dark matter halo scales in the three models. The dark matter mass is maintained constant as $k=10^{4}M$, while the halo characteristic radius $h$ varies from $10^{7}M\sim 10^{10}M$. The several dark matter models produce distinguishable features in gravitational waveforms for a relatively small dark matter halo scale, and the results reveal a convergence pattern as the halo characteristic radius increases. When $h=10^{7}M$, the gravitational waveforms obtained in three dark matter models exhibit noticeable deviations from the Schwarzschild reference baseline, reflecting the significant influence of the dark matter halo on the spacetime geometry and particle orbits. These deviations are evident in waveform characteristics: the signal duration extends to approximately $1200M$ in an orbital time period (compared to roughly $950M$ for Schwarzschild results), and the the oscillations exhibit noticeably distinct characteristics. As the halo characteristic radius expands to $h=10^{8}M$, these deviations diminish substantially, with the waveforms moving closer to the Schwarzschild case. The signal duration contracts toward the Schwarzschild results, and the oscillation behaviors become more similar. This convergence process continues as $h$ increases to $h=10^{9}M$, at which point the waveforms of all dark matter models become nearly indistinguishable from the Schwarzschild reference baseline in terms of signal duration, oscillation behaviors and amplitude structure. When $h>10^{9}M$, the periodic orbits of the dark matter halo models completely coincide with those of the Schwarzschild black hole, and the corresponding gravitational waveforms are identical as well. For this reason, the results for such cases are not presented separately. This convergence behavior directly parallels the halo effects observed in the precession angle analysis (Figs. 8 and 9) and orbital trajectories (Fig. 10). The physical interpretation remains consistent: when the dark matter halo characteristic radius becomes sufficiently large while maintaining dark matter total mass as constant, the dark matter halo is diluted and the local gravitational field gradually approaches that of an isolated Schwarzschild black hole. Thereby reducing the distinguishable signatures of different dark matter density profiles in the gravitational waves.
@@
To provide a more comprehensive view of dark matter halo’s scale effects across different orbital configurations, Fig. 19 presents gravitational waveform comparisons for varying NFW halo scales using five orbital configurations: (110)(1~1~0), (120)(1~2~0), (130)(1~3~0), (211)(2~1~1), and (221)(2~2~1). This figure examines gravitational waveforms for four dark matter scale values: $h=10^{7}M$, $h=10^{8}M$, $h=10^{9}M$ and $h=10^{10}M$, with the Schwarzschild case serving as the reference baseline. These results not only reveal how different orbital configurations (zwv)(z~w~v) modulate the gravitational waveforms (consistent with findings in previous sections), but also indicate the clear trend: larger dark matter halo scales lead to weaker effects on the gravitational potential, with waveforms approaching the Schwarzschild black hole case.
@@ V Conclusion
The effects of dark matter halos on spacetime geometry, particle orbits and gravitational waves can be incorporated into two parameters: the dark matter mass $k$ and halo characteristic radius $h$. Our analysis of the effective potential reveals that the existence of dark matter halos reduces the extrema of the effective potential. This reduction becomes more obvious as the dark matter mass $k$ increases, directly affecting the MBO and ISCO of particles. The presence of dark matter halos increases the angular momentum for both MBO and ISCO, with the NFW and Beta models showing nearly identical behavior, while the Moore model exhibits lower angular momentum values.
@@
To comprehensively explore periodic orbits, we examine the effects of both dark matter mass $k$ and halo characteristic radius $h$ under the condition of selected angular momentum $L=L_{\text{ISCO}}+\varepsilon(L_{\text{MBO}}-L_{\text{ISCO}})$ and the selected energy $E$ (in Appendix B). For the selected angular momentum condition, it is demonstrated that for a given precession angle $q$, larger dark matter masses require lower orbital energies $E$. The NFW and Beta models require substantially lower energies $E_{(z~w~v)}$ than the Moore model for the same precession angle and orbital configurations (zwv)(z~w~v). This energy difference directly affects the orbital shapes, increasing dark matter mass stretches the apoapsis of periodic orbits, which leads to larger orbital dimensions. For a fixed precession angle $q$, larger dark matter halo scales require higher orbital energy. At smaller scales, periodic orbital shapes differ from the Schwarzschild case. As the halo scale increases, the orbits gradually converge toward the Schwarzschild trajectories, closely approaching them at large scales. This behavior reflects the dilution of gravitational effects: when the dark matter mass remains constant while the halo scale increases, the gravitational binding weakens in the regions. Particles must then possess higher energies (closer to the asymptotic value $\lim_{r\to\infty}V_{\text{eff}}=1$ without gravitation) to maintain the same precession characteristics. This scale-dependent effect is clearly shown and explained with five different periodic orbit configurations (zwv)(z~w~v), where orbits gradually approach the Schwarzschild case as $h$ increases. In particular, when the dark matter halo scale is $h\sim 10^{10}M$, the periodic orbits of all three dark matter models become indistinguishable from the Schwarzschild reference, exhibiting the dilution limit of dark matter halo effects. Furthermore, analysis of periodic orbits with fixed $E$ yields conclusions consistent with those obtained for fixed $L$.
@@
To explore the observational signatures of these periodic orbits, we analyze an extreme-mass-ratio inspiral (EMRI) system consisting of a test object with a typical mass ratio and luminosity distance. The gravitational wave analysis shows strong correspondence between the waveforms and orbital properties. Larger dark matter mass leads to more pronounced differences in generated gravitational waves between periodic orbits around SMBH in dark matter halos and those of Schwarzschild black holes, enhancing the distinguishability among three dark matter halo models. The waveforms exhibit stretching trends, particularly an increase in the duration of each complete orbital period, which stems directly from the modification of the gravitational potential by the dark matter halo. In contrast to the dark matter mass effects, a larger dark matter halo scale $h$ results in a more diffuse mass distribution, weakening its effect on the gravitational potential. The gravitational waveforms generated by periodic orbits exhibit significant differences between various dark matter halo models under a relatively small halo scale, which can be used to constrain dark matter models with future observations. The gravitational waveforms gradually approach those for Schwarzschild black holes when dark matter halo scales become sufficiently large. Particularly, when the characteristic radius of the dark matter halo is $h\sim 10^{10}M$, the periodic orbits and their gravitational waveforms converge precisely with those of the Schwarzschild black hole. Furthermore, examining the waveform structure reveals that the orbital configurations (zwv)(z~w~v), as fundamental quantities characterizing the geometric shapes of periodic motion, directly determine the spatiotemporal features (the zoom-whirl stages) of the gravitational wave signals.
@@
Throughout our analysis, we observe that the NFW and Beta models produce nearly indistinguishable results for both orbital characteristics and gravitational wave signals for all parameter ranges examined in this study. This accordance holds for precession angles, periodic orbit configurations, and gravitational waveforms, suggesting that these two density profiles have very similar effects on particle orbits and gravitational radiation in the parameter ranges relevant to galactic dark matter halos. The Moore model, while following similar qualitative trends, generates different quantitative results compared with Beta and NFW models, which become more obvious at larger dark matter masses.
@@
In summary, through an systematic analysis of dark matter halo models, this study clarifies how dark matter halo parameters influence particle orbits around black holes and gravitational wave radiation. The complementary effects of the dark matter mass and its characteristic radius offer a pathway to understand the complex interplay between dark matter halo and SMBH in galactic center, as well as provide theoretical predictions for gravitational waveforms that can be compared with observations. These findings may provide useful constraints for interpreting gravitational wave signals from EMRIs detected by future space-based observatories, such as LISA, TianQin, and Taiji.